# Adaptive Coordinators and Prompts on Heterogeneous Graphs for Cross-Domain Recommendations

### v1

@@ Abstract.
In the online digital world, users frequently engage with diverse items across multiple domains (e.g., e-commerce platforms, streaming services, and social media networks), forming complex heterogeneous interaction graphs. Leveraging this multi-domain information can undoubtedly enhance the performance of recommendation systems by providing more comprehensive user insights and alleviating data sparsity in individual domains. However, integrating multi-domain knowledge for the cross-domain recommendation is very hard due to inherent disparities in user behavior and item characteristics and the risk of negative transfer, where irrelevant or conflicting information from the source domains adversely impacts the target domain’s performance. To address these challenges, we offer HAGO, a novel framework with Heterogeneous Adaptive Graph coOrdinators, which dynamically integrate multi-domain graphs into a cohesive structure by adaptively adjusting the connections between coordinators and multi-domain graph nodes, thereby enhancing beneficial inter-domain interactions while mitigating negative transfer effects. Additionally, we develop a universal multi-domain graph pre-training strategy alongside HAGO to collaboratively learn high-quality node representations across domains. To effectively transfer the learned multi-domain knowledge to the target domain, we design an effective graph prompting method, which incorporates pre-trained embeddings with learnable prompts for the recommendation task. Our framework is compatible with various graph-based models and pre-training techniques, demonstrating broad applicability and effectiveness. Further experimental results show that our solutions outperform state-of-the-art methods in multi-domain recommendation scenarios and highlight their potential for real-world applications. We will open our source code [^1] This paper is currently under review in a conference, we have submitted the source code to the venue. The source link contains the venue information so we will release it after acceptance..
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
In real-world applications, users often interact with various items and other users (Li et al., 2024b) across multiple domains within the platform, such as different product categories in e-commerce or distinct content types in streaming services, forming complex heterogeneous interaction graphs. However, the distribution of data across these domains is often uneven. Some domains benefit from abundant user interaction data, while others suffer from data sparsity. How to effectively utilize the abundant interaction data from well-established source domains to enhance recommendation performance in the target domain that faces data sparsity or holds greater strategic importance to the platform (a.k.a cross-domain recommendation) has become an important area in academia and enterprise.
@@
In recent years, various model-centric methods have been proposed to handle this problem yet still leave a huge gap for further improvement. As illustrated in Figure 1(a), one kind of solution is to build some shared-bottom architectures by utilizing shared representations (Singh and Gordon, 2008; Hu et al., 2018) or representation fusion (Zhu et al., 2019; Gao et al., 2019) to represent and transfer information across different domains. The other direction, as shown in Figure 1(b), is to use some mapping functions or transfer layers to map user or item representations from the source domains to the target domain for information enhancement (Man et al., 2017; Zhao et al., 2020; Kang et al., 2019; Guo et al., 2023). Aligned with these two kinds of model-centric frameworks, they also introduce some advanced training skills such as contrastive learning, to promote cross-domain transfer learning further (Li et al., 2024a; Hao et al., 2024; Xie et al., 2022; Liu et al., 2024). Although these methods have achieved notable success, some later theoretical studies (Cao et al., 2023) have found that the performance ceiling of model-centric methods is strictly limited by the intrinsic discrepancy among various domains, indicating that we might not further improve these methods unless we can find some data-level approaches to narrow down the natural gap among various graph domains (as presented in Figure 1(c)).
@@
However, achieving this goal is never easy because we need to overcome at least three tough challenges: Challenge 1, designing a hand-crafted strategy to align different graph domains is far-fetched and tough. Interaction graph data from different domains may exhibit entirely distinct structural patterns and various types of item nodes, resulting in diverse topologies. This makes manually designed strategies far from effective and general. To this end, we need some trainable approaches to learn the potential latent data operation strategy to align these domains yet preserve their important knowledge. Challenge 2, how to integrate and learn the rich knowledge from multiple source domains is tough. The complexity and diversity of these graphs not only enrich the underlying knowledge but also significantly increase the difficulty for models to capture common patterns across multiple domains. Consequently, effectively learning useful knowledge from multi-domain source graphs presents an intricate challenge, as it is crucial for enabling models to capture cross-domain correlations among entities better. Challenge 3, how to transfer the preserved knowledge to the target is still an open problem, and currently often undergoes serious negative transfer issues. Source domains encompass a vast array of information, much of which may not be pertinent or valuable to the target domain. Indiscriminately transferring knowledge from source domains to the target domain may adversely affect the model’s performance, potentially resulting in a decline in recommendation accuracy.
@@
Inspired by the recent success of graph prompting in its powerful data operation capability (Zhao et al., 2024; Sun et al., 2023a, c), we go beyond the previous model-centric paradigm and hope to bridge gaps between diverse graph datasets in a data-centric perspective. We offer HAGO framework, a data-level framework with Heterogeneous Adaptive Graph coOrdinators to address the aforementioned challenges. Specifically, to address Challenge 1, we revisit the theoretical basis of graph prompts in data manipulations, and then we propose graph coordinators to integrate multi-domain graphs in the pre-training stage, which shares a similar theoretical basis as graph prompts. In this way, we can get rid of hand-crafted strategies and learn a potential way to unify various source domains into a unified structure, enabling collaborative multi-domain pre-training to facilitate representation alignment among them. Notably, our graph coordinators are designed to be heterogeneous, using user/item coordinators to connect user/item nodes in graphs, thus preventing negative transfer caused by information confusion between different types of nodes. To address Challenge 2, we give the details of a multi-domain pre-training framework with our graph coordinator mechanism. This framework is model-agnostic and can be easily used in various pre-training strategies and backbones. To address Challenge 3, the coordinators in HAGO connect the nodes in corresponding graphs by edges with adaptive weights, dynamically regulating knowledge transfer both within and across domains. Moreover, we design an efficient graph prompting method that integrates the pre-trained embeddings and prompts for recommendation tasks. We conduct extensive experiments on two real-world platforms with seven different domains, results from which further demonstrate the effectiveness of our method. In summary, the main contributions of our work can be summarized as follows:
@@ 2. Background
In this section, we introduce our motivation by comparing two perspectives of cross-domain recommendation. Then, the formal objective of this paper will be given. We give a more comprehensive review of existing related work in Appendix A.1.
@@
Cross-domain Recommendation: A Model-centric Perspective. A mainstream approach for cross-domain recommendation is to employ a shared-bottom architectures (Singh and Gordon, 2008; Hu et al., 2018; Zhu et al., 2019; Gao et al., 2019; Zhu et al., 2024) and mapping functions to project knowledge from source domains to the target (Man et al., 2017; Zhao et al., 2020; Kang et al., 2019; Guo et al., 2023). These approaches are inherently model-centric. Although much progress achieved, some researchers recently have found that model-centric approaches have an unbreakable ceiling caused by the intrinsic gap of the cross-domain data (Cao et al., 2023). That means no matter how better the models are, their upper bound performance is always limited by the data natural gap. Thus, it is imperative to return to the intrinsic nature of the problem and consider how to further break up the ceiling. Intuitively, if we can find some way to narrow the data gap among various domains, we will have a higher performance improvement margin for existing models. However, traditional data dealing approaches heavily rely on manpower and it is far from effective and efficient to manually design data operation strategies for this vision.
Figure 2. Illustration of the proposed HAGO, using two source domains and a target domain as examples. There is partial overlap among user nodes across different domains. For the sake of clarity in the diagram, we omit the overlapping user nodes that are present in all domains. HAGO employs heterogeneous adaptive graph coordinators to organize the multi-domain pre-training collaboratively, and then leverages graph prompts to transfer the knowledge gained from pre-training to the target domain, thereby enhancing downstream recommendation tasks.
@@
Our Motivation: An Attempt of Data-centric Perspective. With the above discussion, this paper tries to explore an enhanced direction, which aims to achieve representation alignment and effective knowledge transfer from a data-centric perspective. Unlike other linear-structured data such as images and text, graph data is highly complex and possesses diverse non-linear structural properties. The nodes in a graph are not isolated; they are deeply influenced by their local neighbors. The connectivity among nodes serves as the carrier of information flow within the graph, reflecting the core characteristics of graph data. Consequently, capturing the common patterns of complex and diverse local structures across multiple graphs to align graph data at the data level particularly presents a more challenging problem. It is infeasible to leverage manually designed strategies to heal the dataset gap effectively. Luckily, recent progress on graph prompting has demonstrated the potential effectiveness in manipulating diversified datasets to be unified (Wang et al., 2024). Graph prompting has the potential to bridge the gaps between multi-domain pre-training and knowledge transfer. With this in mind, we aspire to harness the power of graph prompting to amalgamate graphs from various domains into a cohesive structure, fostering alignment and mitigating negative transfer at the data level.
@@
3. Reconcile Multi-Domain Incompatibility via Graph Coordinators and Prompts
Overview: With the above objective, we can see that we should not only try to narrow the gap between the source domains and the target domain, but also need to heal the incompatibility among multiple source domains. To this end, we proposed HAGO as depicted in Figure 2. Specifically, our framework initially employs heterogeneous adaptive graph coordinators to integrate multi-domain graphs into a unified structure and leverages self-supervised learning algorithms for collaborative multi-domain pre-training. Subsequently, graph prompting is strategically implemented to transfer the multi-domain knowledge acquired from pre-training to the target domain, aiming to improve the accuracy of recommendation.
@@ 3.1. A Nutshell of the Backbone Model
where $\boldsymbol{E}_{u}\in\mathbb{R}^{|\mathcal{U}|\times d}$ and $\boldsymbol{E}_{i}\in\mathbb{R}^{|\mathcal{I}|\times d}$ denotes the embedding lookup table of users and items, and $\boldsymbol{o}_{u}\in\mathbb{R}^{|\mathcal{U}|},\boldsymbol{o}_{i}\in\mathbb{R}^{|\mathcal{I}|}$ are the corresponding one-hot vector of user/item IDs. $|\mathcal{U}|$ and $|\mathcal{I}|$ indicate the numbers of users and items, and $d$ means the embedding size.
Propagation Layer: The graph aggregation process of LightGCN can be formalized as:
(2) $\boldsymbol{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}\boldsymbol{e}_{i}^{(l)},$
$\boldsymbol{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{u}\right|}}\boldsymbol{e}_{u}^{(l)},$
@@ 3.2.1. Heterogeneous Graph Coordinators
The interaction data is typically modeled by constructing heterogeneous user-item bipartite graphs, which mainly include two types of nodes: users and items, as well as the user-item edges. Naturally, we introduce a set of user coordinators $\mathcal{C}_{u}^{(k)}=\{c_{u,1}^{(k)},c_{u,2}^{(k)},\cdots,c_{u,n}^{(k)}\}$ and item coordinators $\mathcal{C}_{i}^{(k)}=\{c_{i,1}^{(k)},c_{i,2}^{(k)},\cdots,$ $c_{i,n}^{(k)}\}$ for graph $\mathcal{G}^{(k)}$. The whole set of graph coordinators is $\mathcal{C}=\bigcup_{k}\mathcal{C}_{u}^{(k)}\cup\bigcup_{k}\mathcal{C}_{i}^{(k)}$, and the number of coordinators is $|\mathcal{C}|=2n$. Each graph coordinator connects to the nodes in the corresponding graph through dynamically weighted direct edges, and corresponds to a learnable embedding which can be obtained by retrieving the coordinator embedding table $\boldsymbol{E}_{c}\in\mathbb{R}^{|\mathcal{C}|\times d}$ as the same way in Equation 1.
@@ 3.2.2. Adaptive Connections between Coordinators and Graphs
To enhance the effectiveness of cross-domain knowledge transfer while alleviating the negative transfer phenomenon, we propose an adaptive edge weight learning strategy to dynamically regulate the strength of information transfer between the coordinators and the nodes. Our proposed strategy encompasses both soft/hard cases and can be formalized as follows:
@@ 3.2.3. Interconnections between Coordinators
Through the aforementioned strategy, we establish adaptive linkages between graph coordinators and their respective graphs, empowering these coordinators to bolster individual graph models by regulating the flow of information among the nodes. Our vision, however, transcends this capability, as we aspire to model collaborative connections across multiple domains within our dataset to facilitate multi-domain representation alignment. Consequently, we establish direct cross-domain interconnections between the graph coordinators associated with each domain graph, effectively amalgamating the entire collection of graphs into a cohesive structure. These graph coordinators assume the pivotal role of information exchange hub between graphs, with the connecting edges between coordinators acting as bridges for this communication. Notably, our proposed HAGO framework includes distinct coordinators for users and items, facilitating connections solely between coordinators of different types. This structured relationship can be effectively depicted using an adjacency matrix $\boldsymbol{R}_{C}$:

### v2

# Adaptive Graph Integration for Cross-Domain Recommendation via Heterogeneous Graph Coordinators

@@ Abstract.
In the digital era, users typically interact with diverse items across multiple domains (e.g., e-commerce, streaming platforms, and social networks), generating intricate heterogeneous interaction graphs. Leveraging multi-domain data can improve recommendation systems by enriching user insights and mitigating data sparsity in individual domains. However, integrating such multi-domain knowledge for cross-domain recommendation remains challenging due to inherent disparities in user behavior and item characteristics and the risk of negative transfer, where irrelevant or conflicting information from the source domains adversely impacts the target domain’s performance. To tackle these challenges, we propose HAGO, a novel framework with Heterogeneous Adaptive Graph coOrdinators, which dynamically integrates multi-domain graphs into a cohesive structure. HAGO adaptively adjusts the connections between coordinators and multi-domain graph nodes to enhance beneficial inter-domain interactions while alleviating negative transfer. Furthermore, we introduce a universal multi-domain graph pre-training strategy alongside HAGO to collaboratively learn high-quality node representations across domains. Being compatible with various graph-based models and pre-training techniques, HAGO demonstrates broad applicability and effectiveness. Extensive experiments show that our framework outperforms state-of-the-art methods in cross-domain recommendation scenarios, underscoring its potential for real-world applications. The source code is available at https://github.com/zhy99426/HAGO.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06journalyear: 2025copyright: rightsretainedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italy.booktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italyisbn: 979-8-4007-1592-1/25/07doi: 10.1145/3726302.3729886ccs: Information systems Recommender systems
@@ 1. Introduction
In real-world applications, users often interact with various items (Li et al., 2024b; Meng et al., 2023b) across multiple domains within the platform, such as different product categories in e-commerce or distinct content types in streaming services, forming complex heterogeneous interaction graphs. However, the distribution of data across these domains is often uneven (Meng et al., 2023a). Some domains benefit from abundant user interaction data, while others suffer from data sparsity. How to effectively utilize the abundant interaction data from well-established source domains to enhance recommendation performance in the target domain that faces data sparsity or holds greater strategic importance to the platform (a.k.a, cross-domain recommendation) has become an important area in academia and industry.
@@
In recent years, various model-centric methods have been proposed to address this problem, yet significant room for improvement remains. As illustrated in Figure 1(a), one kind of solution is to build some shared-bottom architectures by utilizing shared representations (Singh and Gordon, 2008; Hu et al., 2018) or representation fusion (Zhu et al., 2019; Gao et al., 2019) to represent and transfer information across different domains. The other direction, as shown in Figure 1(b), is to use some mapping functions or transfer layers to map user or item representations from the source domains to the target domain for information enhancement (Man et al., 2017; Zhao et al., 2020; Kang et al., 2019; Guo et al., 2023). Aligned with these two kinds of model-centric frameworks, they also introduce some advanced training skills such as contrastive learning, to promote cross-domain transfer learning further (Li et al., 2024a; Hao et al., 2024; Xie et al., 2022; Liu et al., 2024). Although these methods have achieved notable success, some recent theoretical studies (Cao et al., 2023) have found that the performance ceiling of model-centric methods is strictly limited by the intrinsic discrepancy among various domains, indicating that we might not further improve these methods unless we can find some data-level approaches to narrow down the natural gap among various graph domains (as presented in Figure 1(c)).
@@
However, achieving this goal is never easy because we need to overcome at least three tough challenges: Challenge 1, designing a hand-crafted strategy to align different graph domains is far-fetched and tough. Interaction graph data from different domains may exhibit entirely distinct structural patterns and various types of item nodes, resulting in diverse topologies. This makes manually designed strategies far from effective and general. To this end, we need some trainable approaches to learn the potential latent data operation strategy to align these domains while preserving their important knowledge. Challenge 2, how to integrate and learn the rich knowledge from multiple source domains is tough. The complexity and diversity of these graphs not only enrich the underlying knowledge but also significantly increase the difficulty for models to capture common patterns across multiple domains. Consequently, effectively learning useful knowledge from multi-domain source graphs presents an intricate challenge, as it is crucial for enabling models to capture cross-domain correlations among entities better. Challenge 3, how to transfer the preserved knowledge to the target is still an open problem, and currently often undergoes serious negative transfer issues. Source domains encompass a vast array of information, much of which may not be pertinent or valuable to the target domain. Indiscriminately transferring knowledge from source domains to the target domain may adversely affect the model’s performance, potentially resulting in an accuracy decline.
@@
Inspired by the recent success of graph prompting in its powerful data operation capability (Zhao et al., 2024; Sun et al., 2023a, b), we go beyond the previous model-centric paradigm and hope to bridge gaps between diverse graph datasets in a data-centric perspective. We offer HAGO framework, a data-level framework with Heterogeneous Adaptive Graph coOrdinators to address the above challenges. Specifically, to address Challenge 1, we revisit the theoretical basis of graph prompts in data manipulations, and then we propose heterogeneous graph coordinators to integrate multi-domain graphs in the pre-training stage, which shares a similar theoretical basis as graph prompts. In this way, we can get rid of hand-crafted strategies and learn a potential way to unify various source domains into a unified structure, enabling collaborative multi-domain pre-training to facilitate representation alignment among them. Notably, our graph coordinators are designed to be heterogeneous, using user/item coordinators to connect user/item nodes in graphs, thus preventing negative transfer caused by information confusion between different types of nodes. To address Challenge 2, we present the details of a multi-domain pre-training framework with our graph coordinator mechanism. This framework is model-agnostic and can be easily used in various pre-training strategies and backbones. To address Challenge 3, the coordinators in HAGO connect the nodes in corresponding graphs by edges with adaptive weights, dynamically regulating knowledge transfer both within and across domains. Moreover, we design an efficient graph prompting method that integrates the pre-trained embeddings and prompts for recommendation tasks. We conduct extensive experiments on two real-world platforms with seven different domains, results from which further demonstrate the effectiveness of our method. In summary, the main contributions of our work can be summarized as follows:
@@ 2. Background
In this section, we introduce our motivation by comparing two perspectives of cross-domain recommendation. Then, the formal objective of this paper will be given.
@@
Cross-domain Recommendation: A Model-centric Perspective. A mainstream approach for cross-domain recommendation is to employ a shared-bottom architecture (Singh and Gordon, 2008; Hu et al., 2018; Zhu et al., 2019; Gao et al., 2019; Zhu et al., 2024) and mapping functions to project knowledge from source domains to the target (Man et al., 2017; Zhao et al., 2020; Kang et al., 2019; Guo et al., 2023). These approaches are inherently model-centric. Despite significant progress, recent studies have revealed that model-centric approaches have an unbreakable ceiling caused by the intrinsic gap of the cross-domain data (Cao et al., 2023). That means no matter how good the models are, their upper bound performance is always limited by the natural data gap. Thus, it is imperative to revisit the intrinsic nature of the problem and explore ways to further break up the ceiling. Intuitively, narrowing the data gap among various domains could provide a higher performance improvement margin for existing models. However, traditional data dealing approaches heavily rely on manpower, making it inefficient and ineffective to manually design data operation strategies for this vision.
@@
Our Motivation: An Attempt of Data-centric Perspective. With the above discussion, this paper tries to explore an enhanced direction, which aims to achieve representation alignment and effective knowledge transfer from a data-centric perspective. Unlike other linear-structured data such as images and text, graph data is highly complex and exhibits diverse non-linear structural properties. Nodes in a graph are not isolated; they are deeply influenced by their local neighbors. The connectivity among nodes serves as the carrier of information flow within the graph, reflecting the core characteristics of graph data. Consequently, capturing the common patterns (Zhang et al., 2024) of complex and diverse local structures across multiple graphs to align graph data at the data level particularly presents a more challenging problem. It is infeasible to leverage manually designed strategies to heal the dataset gap effectively. Fortunately, recent progress on graph prompting has demonstrated the potential effectiveness in manipulating diversified datasets to be unified (Sun et al., 2023a; Wang et al., 2024). Graph prompting has the potential to bridge the gaps between multi-domain pre-training and knowledge transfer. Inspired by this, we aspire to harness the power of graph prompting to amalgamate graphs from various domains into a cohesive structure, fostering alignment and mitigating negative transfer at the data level.
@@
3. Method
Figure 2. Illustration of the proposed HAGO, using two source domains and a target domain as examples. There is partial overlap among user nodes across different domains. For clarity, we omit the overlapping user nodes that are present in all domains. HAGO employs heterogeneous adaptive graph coordinators to organize the multi-domain pre-training collaboratively, and then leverages graph prompts to transfer the knowledge gained from pre-training to the target domain, thereby enhancing downstream recommendation tasks.
Overview: With the above objective, we can see that we should not only try to narrow the gap between the source domains and the target domain, but also need to heal the incompatibility among multiple source domains. To this end, the proposed HAGO is depicted in Figure 2. Specifically, our framework initially employs heterogeneous adaptive graph coordinators to integrate multi-domain graphs into a unified structure and leverages self-supervised learning algorithms for collaborative multi-domain pre-training. Subsequently, graph prompting is strategically implemented to transfer the multi-domain knowledge acquired from pre-training to the target domain, aiming to improve the accuracy of recommendation.
@@ 3.1. A Nutshell of the Backbone Model
where $\boldsymbol{E}_{u}\in\mathbb{R}^{|\mathcal{U}|\times d}$ and $\boldsymbol{E}_{i}\in\mathbb{R}^{|\mathcal{I}|\times d}$ denote the user/item embedding lookup tables, and $\boldsymbol{o}_{u}\in\mathbb{R}^{|\mathcal{U}|},\boldsymbol{o}_{i}\in\mathbb{R}^{|\mathcal{I}|}$ are the corresponding one-hot vector of user/item IDs. $|\mathcal{U}|$ and $|\mathcal{I}|$ indicate the numbers of users and items, and $d$ means the embedding size.
Propagation Layer: The graph aggregation process in LightGCN can be formalized as:
(2) $\boldsymbol{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}\boldsymbol{e}_{i}^{(l)},\boldsymbol{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{u}\right|}}\boldsymbol{e}_{u}^{(l)},$
@@ 3.2. Heterogeneous Adaptive Graph Coordinator
To enhance the interconnections between multi-domains while alleviating the negative transfer phenomenon, we design effective heterogeneous adaptive graph coordinators to organize cross-domain knowledge transfer dynamically.
@@ 3.2.1. Heterogeneous Graph Coordinators
The interaction data is typically modeled by constructing heterogeneous user-item bipartite graphs, which mainly include two types of nodes: users and items, as well as the user-item edges. Naturally, we introduce a set of user coordinators $\mathcal{C}_{u}^{(k)}=\{c_{u,1}^{(k)},c_{u,2}^{(k)},\cdots,c_{u,n}^{(k)}\}$ and item coordinators $\mathcal{C}_{i}^{(k)}=\{c_{i,1}^{(k)},c_{i,2}^{(k)},\cdots,$ $c_{i,n}^{(k)}\}$ for graph $\mathcal{G}^{(k)}$. Here, user coordinators are restricted to connecting only with user nodes, while item coordinators are restricted to connecting only with item nodes. The hyperparameter $n$ determines the number of coordinators for each type, meaning we introduce $2n$ coordinators including $n$ user coordinators and $n$ item coordinators. Each graph coordinator connects to the nodes in the corresponding graph through dynamically weighted direct edges, and corresponds to a learnable embedding which can be obtained by retrieving the coordinator embedding table $\boldsymbol{E}_{c}\in\mathbb{R}^{|\mathcal{C}|\times d}$ as the same way in Equation 1.
@@ 3.2.2. Adaptive Connections between Coordinators and Graphs
To enhance the effectiveness of cross-domain knowledge transfer while alleviating the negative transfer phenomenon, we propose an adaptive edge weight learning strategy to dynamically regulate the strength of information transfer between the coordinators and the nodes. Our proposed strategy includes both soft/hard cases and can be formalized as follows:
@@ 3.2.3. Interconnections between Coordinators
Through the above strategy, we establish adaptive linkages between graph coordinators and their respective graphs, empowering these coordinators to bolster individual graph models by regulating the flow of information among the nodes. Our vision extends further, aiming to model collaborative connections across multiple domains to facilitate multi-domain representation alignment. Consequently, we establish direct cross-domain interconnections between the graph coordinators associated with each domain graph, effectively amalgamating the entire collection of graphs into a cohesive structure. These graph coordinators assume the pivotal role of information exchange hub between graphs, with the connecting edges between coordinators acting as bridges for this communication. Notably, our proposed HAGO framework includes distinct coordinators for users and items, facilitating connections only between coordinators of different types. In other words, user-user and item-item coordinator connections are not included. This design is based on the intuition that, even with the introduction of graph coordinators, the cohesive multi-domain graph should retain the structure of a heterogeneous user-item bipartite graph. This structured relationship can be depicted using an adjacency matrix $\boldsymbol{R}_{C}$: